A box contain 3 dimes, 3 nickels and 1 penny .3 coins are being selected at random without replacement .(Where 1 Dime=10 cent, 1 Nickle=5 cent and 1 penny=1 cent)

1
Expert's answer
2022-03-02T09:32:56-0500

Solution:

3 dimes, 3 nickels and 1 penny

Total = 3+3+1=7

To find the probability that 3 coins selected are dimes.

Required probability "=\dfrac{3}{7}"